OpenStudy (anonymous):

Give that log 4=0.6061, log 9=0.9542, and log 12=1.0792 evaluate each logarithmn Problem: log 108,000

OpenStudy (campbell_st):

108000 = 108 x 1000 = 9 x 12 x 1000 and 1000 = 10^3 and I'd assume you are looking at base 10 logs.... hope it helps

OpenStudy (anonymous):

So, you would multiple log9 times log 12? Then plus 10^3?

OpenStudy (anonymous):

\[\huge \log (108000) = \log(9\times 12 \times 10^3)\]\[\huge\implies \log 9 + \log 12 + \log 10^3\]

OpenStudy (campbell_st):

and it end up as \[\log(9) + \log(12) + 3\log(10) = \log(9) + \log(12) + 3\]